OpenStudy (bossimbacon):

The two-way table shows the number of men and women who voted for Candidate A or B. What percent of people who voted for Candidate A were women? Express your answer to the nearest whole percent. % Men Women Candidate A 1253 1921 Candidate B 1564 1723 @cj49

OpenStudy (michele_laino):

by definition, the requested percentage is: \[\frac{{1921}}{{1921 + 1253}} \times 100 = ...?\]

OpenStudy (bossimbacon):

idk

OpenStudy (michele_laino):

what is : \[1921 + 1253 = ...?\]

OpenStudy (bossimbacon):

3174

OpenStudy (michele_laino):

ok! Now, what is: \[\frac{{1921}}{{3174}} = ...?\]

OpenStudy (bossimbacon):

is it dividing?

OpenStudy (michele_laino):

yes!

OpenStudy (bossimbacon):

0.60

OpenStudy (bossimbacon):

0.605

OpenStudy (michele_laino):

ok! Now what is: 0.605*100=...?

OpenStudy (bossimbacon):

60.5

OpenStudy (michele_laino):

ok! Now we have to round off 60.5 to the nearest unit

OpenStudy (bossimbacon):

so just 60.?

OpenStudy (michele_laino):

better is 61%

OpenStudy (bossimbacon):

ohhhhhh thank you!

OpenStudy (michele_laino):

thank you!